OpenStudy (tayfishb):

find a point c satisfying the conclusion of the Mean Value Theorem for the given function and interval y=cosx-sinx [0,2pi]

OpenStudy (tayfishb):

\[f \prime c=\frac{ f(b)-f(a) }{ b-a }\]

OpenStudy (tayfishb):

\[f \prime x=-sinx-cosx\]

OpenStudy (tayfishb):

\[\frac{ (-\cos2\pi -\sin2\pi )-(\cos 0-\sin 0) }{2\pi -0 }\]

OpenStudy (tayfishb):

\[\frac{ (1-0)-(1-0) }{ 2\pi -0}=\frac{ 0 }{ 2 \pi }\]

OpenStudy (tayfishb):

\[0=-sinx-cosx\]

OpenStudy (tayfishb):

\[sinx=-cosx\]

OpenStudy (tayfishb):

now what?????

OpenStudy (tayfishb):

@mathmale @TheSmartOne

Nnesha (nnesha):

divide both sides by -cosx

OpenStudy (tayfishb):

\[-tanx=1???\]

Nnesha (nnesha):

yes solve for tanx and then use the unit circle to find where tanx = -1

Nnesha (nnesha):

\[\frac{\color{Red}{-} \cos 2\pi - \sin 2\pi }{ 2\pi -0 }\] `-` is a typo i guess

OpenStudy (tayfishb):

but arent there two answers for that? second and fourth quadrant?

Nnesha (nnesha):

yes.

OpenStudy (tayfishb):

so \[x=\frac{ \sqrt{2} }{ 2}\]?

OpenStudy (tayfishb):

no wait nevermind i got it

Nnesha (nnesha):

hmm in other words you're solving for x to find an angle you should take invs of trig ratio so x= arctan(-1) sqrt{2}/2 is just the x-coordinate /y-coordinate the tanx tanx=sinx/cosx so basically you have to find at what degree/radian tanx=-1

Nnesha (nnesha):

x-coordinate /y-coordinate NOT the tanx **

OpenStudy (tayfishb):

yeah i got it, its \[\frac{ 7\pi }{ 4 }\]

OpenStudy (tayfishb):

thank you sooo much

Nnesha (nnesha):

what about the other one there are two right ?? b/w [0,2pi]

OpenStudy (tayfishb):

my textbook answers said only the 7pi/4 was correct